OpenStudy (hellomynameiskam):

Solve the equation. - 2/5n = -30 A. n = –75 B. n = –12 C. n = 12 D. n = 75

Mehek (mehek14):

\(\tt{n=\dfrac{-30}{1}*\dfrac{5}{-2}}\)

OpenStudy (hellomynameiskam):

150/-2

Mehek (mehek14):

\(\dfrac{-150}{-2}=\dfrac{150}{2}=?\)

OpenStudy (hellomynameiskam):

75

Mehek (mehek14):

n = 75